# TrojanForge: Adversarial Hardware Trojan Examples with Reinforcement Learning

### v1

@@ Abstract
The Hardware Trojan (HT) problem can be thought of as a continuous game between attackers and defenders, each striving to outsmart the other by leveraging any available means for an advantage. Machine Learning (ML) has recently been key in advancing HT research. Various novel techniques, such as Reinforcement Learning (RL) and Graph Neural Networks (GNNs), have shown HT insertion and detection capabilities. HT insertion with ML techniques, specifically, has seen a spike in research activity due to the shortcomings of conventional HT benchmarks and the inherent human design bias that occurs when we create them. This work continues this innovation by presenting a tool called “TrojanForge”, capable of generating HT adversarial examples that defeat HT detectors; demonstrating the capabilities of GAN-like adversarial tools for automatic HT insertion. We introduce an RL environment where the RL insertion agent interacts with HT detectors in an insertion-detection loop where the agent collects rewards based on its success in bypassing HT detectors. Our results show that this process leads to inserted HTs that evade various HT detectors, achieving high attack success percentages. This tool provides insight into why HT insertion fails in some instances and how we can leverage this knowledge in defense.
@@
1 Introduction
The economics of the silicon supply chain has driven silicon vendors to adopt a globalized fabless business model to curb costs. Most design layouts are sent overseas to be manufactured at semiconductor foundries Bhunia and Tehranipoor (2018). This has introduced a significant increase in revenue for many companies, where the share of the fabless semiconductor firms’ revenue reached 33% of the entire chip industry in 2020 from just 7.6% in 2000 Xing et al. (2023). Nevertheless, this success comes with a cost to chip security. Outsourced designs can be subject to unwanted malicious design modifications such as Hardware Trojans (HT), i.e., malicious embeddings that can be introduced at the chip design or manufacturing stage. An HT has a triggering component that defines the conditions under which the HT is activated, and upon activation, the payload of the HT, which is the malicious action or behavior, can cause a variety of threats that compromise the Confidentiality, Integrity, or Availability of hardware devices, such as Denial of Service (DoS), information leakage, reducing transistor age, and thermal leakage Xue et al. (2020).
To study the impact of (HTs), researchers have primarily used the dataset available at Trust-hub.org Shakya et al. (2017). HT benchmarks are essential assets because they provide opportunities to study the impact of various HTs and the development of HT detection methods. Despite the valuable effort to create these benchmarks for the research community, several associated shortcomings exist. Krieg Krieg (2023) has studied these benchmarks from correctness, maliciousness, stealthiness, and persistence standpoints. He deduced that only three out of 83 studied HT circuits are usable in real-world scenarios. The major issues were pre-synthesis and post-synthesis discrepancies, unsatisfiable trigger conditions, and incorrect original designs. In addition to this, the Trusthub benchmark set is restricted in size and diversity, hindering the ability to train efficient HT detectors that rely on large training datasets. Last but not least, the inherent human bias in HT insertion in Trusthub circuits is another serious problem that undermines the benchmark’s quality. These shortcomings have urged the community to find alternative solutions to these benchmarks in the past few years by developing various automated HT insertion tools Cruz et al. (2018); Gohil et al. (2022a); Sarihi et al. (2022); Gohil et al. (2024).
@@ 1 Introduction
In the Machine Learning (ML) domain, adversarial examples are crafted inputs that are manipulated to mislead machine learning models and force them to return wrong predictions. For instance, McAfee researchers famously tricked a self-driving car into speeding to 85 miles per hour (mph) by placing a 2-inch piece of duct tape on a 35 mph road sign Povolny and Trivedi (2020). This work continues to innovate by introducing a framework that can insert (generate) HT adversarial examples with Reinforcement Learning (RL). We seek inspiration from Generative Adversarial Networks (GANs) to generate new and better HT insertions. We utilize an RL agent that generates HT instances intending to evade HT detectors, i.e., the GAN’s generator component. We propose the “TrojanForge”framework in which an HT inserter tool operates in conjunction with HT detectors (the GAN’s Discriminator) in a loop and tries to collect the maximum positive reward by generating HT instances that are not detectable.
This work’s threat model assumes HT insertion either at design or manufacturing time, i.e., a malicious employee has access to the state-of-the-art reverse-engineering techniques in a foundry, reverse-engineers a design to obtain the netlist, and embeds HTs into this netlist. Similar threat assumptions have been made Gohil et al. (2022a, b). We assume that the adversary uses existing HT detectors or can develop them. Figure 1 shows an overview of our proposed framework where a selection of trigger candidate nets are picked from a netlist and are passed to an RL HT insertion agent. (More details will be explained in Section 3.)
@@
  • We utilize the Jaccard Similarity Index Pandit et al. (2011) to diversify candidate nets for HT insertion. This helps the HT inserting agent to proceed when HTs are hard to insert. The metric also provides insight into designing and defending HT insertion.
@@
2 Background and Related Work
In this section, we discuss HT insertion and detection.
2.1 HT Insertion tools
Initial attempts to gather an HT benchmark were realized in TrustHub, where a list of 96 benchmarks are available to use Shakya et al. (2017); Salmani et al. (2013). Despite the valuable novelty, the dataset suffers from shortcomings such as limited size and diversity Krieg (2023).
@@ 2.1 HT Insertion tools
Various studies have been conducted to address Trusthub’s shortcomings. Cruz et al. Cruz et al. (2018) offers an automated HT generation tool that inserts HTs based on entered parameters from the user. They include the total number of trigger nets, the number of rare trigger nets, the number of inserted HT instances, and the type of payload. The location of a payload is selected randomly. Sarihi et al. Sarihi et al. (2022) introduced an RL-based tool in which an agent explores circuits and takes five different positional actions to insert HT instances in a way that maximizes the sum of collected rewards. The agent receives rewards according to the number of engaged circuit inputs in the HT activation test vector. Gohil et al. Gohil et al. (2022a) proposed another HT insertion RL tool, ATTRITION. ATTRITION sets a signal probability threshold for trigger candidates, and the RL agent is rewarded proportional to the size of “compatible” trigger rare nets it finds. Compatible rare nets are a set of rare nets that can be activated with a single test vector. Nozawa et al. Nozawa et al. (2021) degrade the performance of ML-based HT detectors by introducing adversarial HT examples. The approach changes the structure of the netlist while preserving its functionality. These changes tamper with the feature values of the circuit and, subsequently, cause a circuit misclassification by the detector. The authors use Trusthub HTs to showcase the success of their method. Sarihi et al. Sarihi et al. (2024a) uses ABC, an open-source synthesis tool, to restructure graphs while preserving their functionality. The goal is to introduce a new benchmark with unseen data points. The authors demonstrate that current state-of-the-art HT detectors cannot detect most of the proposed benchmarks. Gohil et al. Gohil et al. (2024) introduced AttackGNN, a tool that integrates ABC with an RL agent to generate adversarial examples that mislead GNNs (graph neural networks) which are used in four hardware security applications. The RL agent uses a combination of graph restructuring actions to achieve the highest attack rates. Saravanan Saravanan et al. (2023) et al. propose a rare net selection technique that has two main parts. First, the authors use the structural information of circuits such as the number of unique primary inputs and gates that form the input logic cone of a rare net to analyze each rare net. Second, random sampling with a tree-based search is used to uncover overlapped trigger combinations and score them based on their quality.
2.2 HT Detection Tools
Our proposed framework needs a discriminator (HT detector) tool. Here, we review HT detection.
@@ 2.2 HT Detection Tools
TGRL Pan and Mishra (2021) is an RL framework that generates test vectors by tweaking the input bit patterns to activate a set of rare nets. The algorithm shows higher trigger coverage compared to MERO; however, neither the test vectors are released, nor the vectors are tested on real HT benchmarks. DETERRENT Gohil et al. (2022b) also employs RL to find the largest set of rare nets that are compatible Gohil et al. (2022b). The authors discuss how to prune some actions for a faster convergence; however, they do not consider rare net pruning. Sarihi et al. Sarihi et al. (2023) introduces three detection proposals based on SCOAP (Sandia Controllability and Observability Analysis Program parameters) and signal switching activity to generate test vectors for ISCAS-85 benchmarks. The authors combine all the test vectors Sarihi et al. (2024b) and demonstrate their efficacy against HT implementations. Hasegawa et al. Hasegawa et al. (2022) introduces a robust ML HT detector called R-HTDetector trained with 51 features extracted from TrustHub benchmarks. The detector aims to increase resiliency against the adversarial gate modification attacks proposed in Nozawa et al. (2021).
@@ 3 TrojanForge
In this section, we describe our proposed framework, TrojanForge, consisting of the tools that conduct a GAN-like HT insertion. The first step in the process is to obtain the gate-level netlist of the design. Next, a Python script converts the netlist to an equivalent acyclic-directed graph representation. A graph consists of a set of vertices and edges $G=(V,E)$, where $V$ is a set of vertices (gates), and $E$ is a set of edges (nets), such that $E\subseteq(V\times V$). Before the RL agent starts processing the graph for HT insertion, a set of candidate nets will be used to build the HT triggers. We will use the terms nets and edges interchangeably in this work. We also assume that our HTs are combinational circuits with an XOR payload similar to Cruz et al. (2018); Gohil et al. (2022a); Sarihi et al. (2022).
@@ 3.1 Rare Net pruning
Since each design consists of thousands of nets, choosing candidate nets for HT trigger circuitry is costly. In this section, we propose a heuristic to reduce the search space. This process starts with a set of 50,000 random test vectors simulated on each netlist, and we record the resulting internal net values. A filter removes candidate nets with a switching probability greater than $th_{sw}$, similar to Cruz et al. (2018); Gohil et al. (2022a). The remaining nets will be referred to as rare nets denoted as set $T$. This initial pruning step is essentially a probability-based filtering. Next, we execute two sequential steps: Functional Pruning described in Section 3.1.1 and Structural Pruning in Section 3.1.2. Both steps aim to further narrow down the candidate set, $T$, reducing the search space for the RL agent and shortening its training time.
@@ 3.1.2 Structural Pruning
Before starting this process, we compute the HT trigger susceptibility parameter Sebt et al. (2018) for the already functionally pruned nets. Equation 2 describes this parameter:
@@
where $HTS$ is the HT trigger susceptibility parameter of $Net_{i}$. $CC0(Net_{i})$ and $CC1(Net_{i})$ are the combinational controllability 00 and $1$ of $Net_{i}$, respectively. $HTS$ stems from the fact that low-switching nets have mainly a high difference between their controllability values. $HTS$ falls within $[0,1)$, and higher values correlate with lower activity on the net. The Structural Pruning process involves deriving a feature vector for each trigger candidate in the graph. Table 1 displays a snapshot of such properties for four nets in $c3540$ after Functional Pruning where Edge denotes the edge id, $RV$ shows the rare value, $SW$ is the switching probability, and $Level$ is the logic level of the net. Also, for each edge, we derive its functionality from its logic input cone. As can be seen from the table, these nets display similar structural behaviors resembling the behavior of a bus in the circuit. While they might not share the exact input logic cone, their functionality shows a high resemblance. Hence, instead of including these nets in the candidate rare net set, we only pick one as a representative net in this pruning process. This step leaves us with a set of nets with unique $HTS$ values upon completion. We denote this set as $S_{pruned}$ where $S_{pruned}\subseteq F_{pruned}$.
@@ 3.2 Adversarial Training
When configured for 5-input HT insertion, the RL agent uses a vector $A$, defined as $A=[a_{0},a_{1},a_{2},a_{3},a_{4}]$, to represent the action space, where each $a_{i}\in[0,\lvert M\rvert-1]$. A state space vector $S$ is the sorted version of the action space, i.e., $S=sort([a_{0},a_{1},a_{2},a_{3},a_{4}])$. If a repeated action (trigger) is selected, it will be replaced by a random one. Next, the payload target net $P$ is selected from all the nets except the ones in $M$. $P$ has a higher logic level than all of the trigger candidates. The payload selection constraints ensure no combinational loops in the final circuit. We use an ATPG algorithm (used in Sarihi et al. (2022)) to ensure that inserted HTs are functional.
@@ 3.3 Special Case: Incompatible Triggers
where $L_{i}$ and $L_{j}$ contain the list of PIs (primary inputs) in the input logic cone of the $i_{th}$ and the $j_{th}$ triggers, respectively. This parameter ranges between $[0,1]$ where higher $JSI$ values indicate more upstream logic cone similarity between triggers, possibly leading to less viable insertions. When the set of trigger candidate nets has a high $JSI$ value, our heuristic approach is to replace some candidate nets with regular nets to lower $JSI$. The set of regular nets is denoted as $R$ where $R\subseteq E-T$. When there are indications of high $JSI$, we set a threshold $JSI_{th}$ and group nets with similar JSI thresholds together. We keep one net from each group with the smallest input logic cone size and replace the remaining nets with regular nets. The replaced nets do not share any input logic cones with the existing set. This process introduces diversity into the trigger choice. We note that this step of the proposed framework emerged from the analysis of our experiments. The $JSI$ metric and our heuristic approach here suggest that the metric might be used as a design goal for a circuit to help protect it from HT insertions.
@@ 4 Experimental Results
In this section, our focus is to demonstrate the capabilities of our GAN-like adversarial HT insertion system and demonstrate how it can evade HT detection. For our experiments, we use an AMD EPYC 7702P 64-core CPU with 512GB of RAM to pre-process graphs, train, and test the RL agents. RL agents are trained with the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017) from the Stable Baselines library Raffin et al. (2021). We target ISCAS-85 benchmark circuits for our study and use NetworkX Hagberg et al. (2008) to process graphs. The RL episode length is set to 25, and with each environment reset, we assign a new payload net to enable the RL agent to experiment with different payloads, thus preventing it from becoming stuck with ones that cannot effectively propagate the HTs’ impact to the circuit outputs. Publicly available test vectors from DETERRENT Gohil et al. (2022b) and three sets of test vectors from Sarihi et al. Sarihi et al. (2023) are used for HT detection. In this study, $M=20$. If the attacker wants to explore higher $M$ values, RL algorithms can deal with large, complex environments.
For our experiments, we use four detectors as the discriminator component of the loop, and we will report our results based on the pairing of our HT insertion with the appropriate discriminator. We note that the nature of our circuit benchmarks is both small and simple and does not reflect modern circuit design. This is a noted shortcoming in both HT detection and insertion space, which we are helping to improve with ML-based tools such as the TrojanForge.

### v3

# TrojanForge: Generating Adversarial Hardware Trojan Examples Using Reinforcement Learning

@@ Abstract
The Hardware Trojan (HT) problem can be thought of as a continuous game between attackers and defenders, each striving to outsmart the other by leveraging any available means for an advantage. Machine Learning (ML) has recently played a key role in advancing HT research. Various novel techniques, such as Reinforcement Learning (RL) and Graph Neural Networks (GNNs), have shown HT insertion and detection capabilities. HT insertion with ML techniques, specifically, has seen a spike in research activity due to the shortcomings of conventional HT benchmarks and the inherent human design bias that occurs when we create them. This work continues this innovation by presenting a tool called “TrojanForge”, capable of generating HT adversarial examples that defeat HT detectors; demonstrating the capabilities of GAN-like adversarial tools for automatic HT insertion. We introduce an RL environment where the RL insertion agent interacts with HT detectors in an insertion-detection loop where the agent collects rewards based on its success in bypassing HT detectors. Our results show that this process helps inserted HTs evade various HT detectors, achieving high attack success percentages. This tool provides insight into why HT insertion fails in some instances and how we can leverage this knowledge in defense.
@@
The economics of the silicon supply chain has driven silicon vendors to adopt a globalized fabless business model to curb costs. Most design layouts are sent overseas to be manufactured at semiconductor foundries [1]. This has introduced a significant increase in revenue for many companies, where the share of the fabless semiconductor firms’ revenue has reached 33% of the entire chip industry in 2020 from just 7.6% in 2000 [2]. Nevertheless, this success comes with a security cost imposed to such chips i.e., outsourced designs can be subject to unwanted malicious design modifications such as Hardware Trojans (HT), i.e., malicious embeddings that can be introduced at the chip design or manufacturing stage. An HT has a triggering component that defines the conditions under which the HT is activated, and upon activation, the payload of the HT, which is the malicious action or behavior, can cause a variety of threats that compromise the Confidentiality, Integrity, or Availability of hardware devices, such as Denial of Service (DoS), information leakage, reducing transistor age, and thermal leakage [3].
To study the impact of (HTs), researchers have primarily used the dataset available at Trust-hub.org [4]. HT benchmarks are essential assets because they provide opportunities to study the impact of various HTs and the development of HT detection methods. Despite the valuable effort to create these benchmarks for the research community, several associated shortcomings exist. Krieg [5] has studied these benchmarks from correctness, maliciousness, stealthiness, and persistence standpoints. He deduced that only three out of 83 studied HT circuits are usable in real-world scenarios. The major issues were pre-synthesis and post-synthesis discrepancies, unsatisfiable trigger conditions, and incorrect original designs. In addition to this, the Trusthub benchmark set is restricted in size and diversity, hindering the ability to train efficient HT detectors that rely on large training datasets. Last but not least, the inherent human bias in HT insertion in Trusthub circuits is another serious problem that undermines the benchmark’s quality. These shortcomings have urged the community to find alternative solutions to these benchmarks in the past few years by developing various automated HT insertion tools [6, 7, 8, 9].
@@
In the Machine Learning (ML) domain, adversarial examples are crafted inputs that are manipulated to mislead machine learning models and force them to return wrong predictions. Following the same paradigm, we seek inspiration from Generative Adversarial Networks (GANs) to generate novel and more challenging HT insertions. Our proposed Reinforcement Learning (RL) agent, resembling the GAN’s generator component, generates HT instances intending to evade HT detectors. We propose the “TrojanForge”framework in which an HT inserter tool operates in conjunction with HT detectors (the GAN’s Discriminator) in a loop and tries to collect the maximum positive reward by generating HT instances that are not detectable. The generated HT instances can be found in [10]. Figure 1 shows an overview of our proposed framework where a selection of trigger candidate nets are picked from a netlist and are passed to an RL HT insertion agent. (More details will be explained in Section 2.)
This work’s threat model assumes HT insertion either at design or manufacturing time, i.e., a malicious employee who has access to the state-of-the-art reverse-engineering techniques in a foundry, reverse-engineers a design to obtain the netlist, and embeds HTs into this netlist. Similar threat assumptions have been made [7, 11]. We assume that the adversary uses existing HT detectors or can develop them.
@@
  • We utilize the Jaccard Similarity Index [12] to diversify candidate nets for HT insertion. This helps the HT inserting agent to proceed when HTs are hard to insert. The metric also provides insight into designing and defending HT insertion.
@@
1 Review of HT Insertion & Detection Literature
Initial attempts to gather an HT benchmark were realized in TrustHub, where a list of 96 benchmarks are available to use [4, 13]. Despite the valuable novelty, the dataset suffers from shortcomings such as limited size and diversity [5]. Various studies have been conducted to address Trusthub’s shortcomings. Cruz et al. [6] offers an automated HT generation tool that inserts HTs based on entered parameters from the user. They include the total number of trigger nets, the number of rare trigger nets, the number of inserted HT instances, and the type of payload. The location of a payload is selected randomly.
@@ 1 Review of HT Insertion & Detection Literature
Sarihi et al. [8] introduced an RL-based tool in which an agent explores circuits and takes five different positional actions to insert HT instances in a way that maximizes the sum of collected rewards. The agent receives rewards according to the number of engaged circuit inputs in the HT activation test vector. Gohil et al. [7] proposed another HT insertion RL tool, ATTRITION. ATTRITION sets a signal probability threshold for trigger candidates, and the RL agent is rewarded proportional to the size of “compatible” trigger rare nets it finds. Compatible rare nets are a set of rare nets that can be activated with a single test vector. Nozawa et al. [14] degrade the performance of ML-based HT detectors by introducing adversarial HT examples. The approach changes the structure of the netlist while preserving its functionality. These changes tamper with the feature values of the circuit and, subsequently, cause a circuit misclassification by the detector. The authors use Trusthub HTs to showcase the success of their method. Sarihi et al. [15] uses ABC, an open-source synthesis tool, to restructure graphs while preserving their functionality. The goal is to introduce a new benchmark with unseen data points. The authors demonstrate that current state-of-the-art HT detectors cannot detect most of the proposed benchmarks. Gohil et al. [9] introduced AttackGNN, a tool that integrates ABC with an RL agent to generate adversarial examples that mislead GNNs (graph neural networks) which are used in four hardware security applications. The RL agent uses a combination of graph restructuring actions to achieve the highest attack rates.
Saravanan [16] et al. propose a rare net selection technique that has two main parts. First, the authors use the structural information of circuits such as the number of unique primary inputs and gates that form the input logic cone of a rare net to analyze each rare net. Second, random sampling with a tree-based search is used to uncover overlapped trigger combinations and score them based on their quality.
@@
As our proposed framework needs a discriminator (HT detector) tool, in the rest of this section, we review HT detection literature. TGRL [17] is an RL framework that generates test vectors by tweaking the input bit patterns to activate a set of rare nets. The algorithm shows higher trigger coverage compared to MERO; however, neither the test vectors are released, nor the vectors are tested on real HT benchmarks. DETERRENT [11] also employs RL to find the largest set of rare nets that are compatible [11]. The authors discuss how to prune some actions for a faster convergence; however, they do not consider rare net pruning. Sarihi et al. [18] introduces three detection proposals based on SCOAP (Sandia Controllability and Observability Analysis Program parameters) and signal switching activity to generate test vectors for ISCAS-85 benchmarks. The authors combine all the test vectors [19] and demonstrate their efficacy against HT implementations. Hasegawa et al. [20] introduces a robust ML HT detector called R-HTDetector trained with 51 features extracted from TrustHub benchmarks. The detector aims to increase resiliency against the adversarial gate modification attacks proposed in [14].
@@ 2 TrojanForge
In this section, we describe our proposed framework, TrojanForge, consisting of the tools that conduct a GAN-like HT insertion. The first step in the process is to obtain the gate-level netlist of the design. Next, a Python script converts the netlist to an equivalent acyclic-directed graph representation. A graph consists of a set of vertices and edges $G=(V,E)$, where $V$ is a set of vertices (gates), and $E$ is a set of edges (nets), such that $E\subseteq(V\times V$). Before the RL agent starts processing the graph for HT insertion, a set of candidate nets will be used to build the HT triggers. We will use the terms nets and edges interchangeably in this work. We also assume that our HTs are combinational circuits with an XOR payload similar to [6, 7, 8].
@@ 2.1 Rare Net Pruning
Since each design consists of thousands of nets, choosing candidate nets for HT trigger circuitry is costly. In this section, we propose a heuristic to reduce the search space. This process starts with a set of 50,000 random test vectors simulated on each netlist, and we record the resulting internal net values. A filter removes candidate nets with a switching probability greater than $th_{sw}$, similar to [6, 7]. The remaining nets will be referred to as rare nets denoted as set $T$. This initial pruning step is essentially a probability-based filtering. Next, we execute two sequential steps: Functional Pruning described in Section 2.1.1 and Structural Pruning in Section 2.1.2. Both steps aim to further narrow down the candidate set, $T$, reducing the search space for the RL agent and shortening its training time.
@@ 2.1.2 Structural Pruning
Before starting this process, we compute the HT trigger susceptibility parameter [21] for the already functionally pruned nets. Equation 2 describes this parameter:
@@
where $HTS$ is the HT trigger susceptibility parameter of $Net_{i}$. $CC0(Net_{i})$ and $CC1(Net_{i})$ are the combinational controllability 00 and $1$ of $Net_{i}$, respectively. $HTS$ stems from the fact that low-switching nets have mainly a high difference between their controllability values. $HTS$ falls within $[0,1)$, and higher values correlate with lower activity on the net. The Structural Pruning process involves deriving a feature vector for each trigger candidate in the graph. Table 1 displays a snapshot of such properties for four nets in $c3540$ after a functional pruning where Edge denotes the edge id, $RV$ shows the rare value, $SW$ is the switching probability, and $Level$ is the logic level of the net. Also, for each edge, we derive its functionality from its logic input cone. As can be seen from the table, these nets display similar structural behaviors resembling the behavior of a bus in the circuit. While they might not share the exact input logic cone, their functionality shows a high resemblance. Hence, instead of including these nets in the candidate rare net set, we only pick one as a representative net in this pruning process. This step leaves us with a set of nets with unique $HTS$ values upon completion. We denote this set as $S_{pruned}$ where $S_{pruned}\subseteq F_{pruned}$.
@@ 2.2 Adversarial Training
When configured for 5-input HT insertion, the RL agent uses a vector $A$, defined as $A=[a_{0},a_{1},a_{2},a_{3},a_{4}]$, to represent the action space, where each $a_{i}\in[0,\lvert M\rvert-1]$. A state space vector $S$ is the sorted version of the action space, i.e., $S=sort([a_{0},a_{1},a_{2},a_{3},a_{4}])$. If a repeated action (trigger) is selected, it will be replaced by a random one. Next, the payload target net $P$ is selected from all the nets except the ones in $M$. $P$ has a higher logic level than all of the trigger candidates. The payload selection constraints ensure no combinational loops in the final circuit. We use an ATPG algorithm (used in [8]) to ensure that inserted HTs are functional.
@@ 2.3 Special Case: Incompatible Triggers
where $L_{i}$ and $L_{j}$ contain the list of PIs (primary inputs) in the input logic cone of the $i_{th}$ and the $j_{th}$ triggers, respectively. The operator $\lvert x\lvert$ returns the number of items in the list $x$. $JSI$ ranges between $[0,1]$ where higher values indicate more upstream logic cone similarity between triggers, possibly leading to less viable insertions. When the set of trigger candidate nets has a high $JSI$ value, our heuristic approach is to replace some candidate nets with regular nets to lower $JSI$. The set of regular nets is denoted as $R$ where $R\subseteq E-T$. When there are indications of high $JSI$, we set a threshold $JSI_{th}$ and group nets with similar JSI thresholds together. We keep one net from each group with the smallest input logic cone size and replace the remaining nets with regular nets. The replaced nets do not share any input logic cones with the existing set. This process introduces diversity into the trigger choice. We note that this step of the proposed framework emerged from the analysis of our experiments. The $JSI$ metric and our heuristic approach here suggest that the metric might be used as a design goal for a circuit to help protect it from HT insertions.
@@ 3 Experimental Results
In this section, our focus is to demonstrate the capabilities of our GAN-like adversarial HT insertion system and demonstrate how it can evade HT detection. For our experiments, we use an AMD EPYC 7702P 64-core CPU with 512GB of RAM to pre-process graphs, train, and test the RL agents. RL agents are trained with the Proximal Policy Optimization (PPO) algorithm [22] from the Stable Baselines library [23]. We target ISCAS-85 [24] benchmark circuits for our study and use NetworkX [25] to process graphs. The RL episode length is set to 25, and with each environment reset, we assign a new payload net to enable the RL agent to experiment with different payloads, thus preventing it from becoming stuck with ones that cannot effectively propagate the HTs’ impact to the circuit outputs. Publicly available test vectors from DETERRENT [11] and three sets of test vectors from Sarihi et al. [18] are used for HT detection. In this study, $M=20$. If the attacker wants to explore higher $M$ values, RL algorithms can deal with large, complex environments. We will report our results based on the pairing of our HT insertion with the appropriate discriminator. We note that the nature of our circuit benchmarks is both small and simple and does not reflect modern circuit design. This is a noted shortcoming in both HT detection and insertion space, which we are helping to improve with ML-based tools such as the TrojanForge.